Assume a ≥ b ≥ c ≥ d If the mean of a , b , c and d, is 8 and the median of
a , b , c and d, is 5 then a + d is
a. 6 b. 11 c. 22 d. 27 e Impossible to Determine

Answer by Theo(13342)

the median is 5 means that (b + c) / 2 = 5

this means that (b + c) = 10 assuming you multiply both sides of that equation by 2.

the mean is 8 means that (a + b + c + d)/4 = 8

substituting for (b + c) = 10, we get:

(a + 10 + d) / 4 = 8

multiply both sides of this equation by 4 to get:

a + d + 10 = 32

subtract 10 from both sides of this equation to get:

a + d = 22

your answer would be selection c (22).